OpenStudy (anonymous):

Which of the following equations has infinitely many real solutions? 5x + 1 = x - 1 5(x - 5) = 5x + 25 5x + 1 = 5x - 1 5(x - 5) = 5x - 25

OpenStudy (anonymous):

Take a look at the last one. \[5(x-5)=5x-25\] Multiply in the parentheses \[5x-25=5x-25\] add 25 \[5x=5x\] divide with 5 \[x=x\] Here you see that x can be anything

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

u were right!

OpenStudy (anonymous):

I hope you understand how I got the answer, that is more important than getting something correct in a test :)

OpenStudy (anonymous):

Yeah thanks, I understand and this was study island